# A Survey on Adversarial Contention Resolution

### v1

@@ 3.2.1. Deterministic $k$-Selection
With regard to upper bounds, De Marco and Kowalski (doi:10.1137/140982763, ) consider the case with a global clock, as discussed above. Subsequently, De Marco et al. (DBLP:conf/icdcs/MarcoKS19, ) examine the power of acknowledgement-based versus non-acknowledgement-based deterministic non-adaptive approaches when $k$ is known; there is neither a global clock, nor is there CD. They give a non-acknowledgement-based algorithm with latency $O(k^{2}\log n)$, and show that this can be improved to $O(k^{2}\log n/\log k)$ with an acknowledgment-based algorithm. These results are nearly tight; the authors also provide lower bounds on latency of $\Omega(k^{2}/\log k)$ for both cases. With regards to throughput, these results by De Marco et al. (DBLP:conf/icdcs/MarcoKS19, ) demonstrate an exponential separation between the case when $k$ is known and $k$ is unknown. For acknowledgement-based algorithms, when $k$ is known, the throughput is $\Omega(1/(\log k\log n))$. In contrast, when $k$ is unknown, De Marco et al. show that throughput is $\Omega(1/(k\log n))$, and the best throughput possible is $O(1/(k\log k))$.
@@ 3.2.2. Deterministic Wakeup
Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) consider a multi-channel scenario with no-CD and no global clock. Here, all stations may access all $b\geq 1$ channels at once. The authors give a non-adaptive, deterministic algorithm for the cases where $k$ is unknown that has latency $O(k\log^{1/b}k\log n)$. When $b=\Omega(\log\log n)$, a faster deterministic algorithm is possible, with delay $O((k/b)\log n\log(b\log n))$, and a lower bound of $((k/b)\log(n/k))$ is shown for deterministic algorithms.
Sticking with a model where there is no-CD and no global clock, De Marco and Kowalski (DEMARCO20171, ; de2013contention, ) consider non-adaptive solutions, with a focus on the time, $s$, until the first station awakens. The authors look at three settings: where $s$ is known and $k$ is unknown, when $k$ is known and $s$ is unknown, and when both $k$ and $s$ are unknown. In the first two settings, the authors give algorithms that achieve a latency of $O(k\log(n/k)+1)$, while their solution for the third scenario has a latency of $O(k\log(n)\log\log(n))$. A lower bound of $\Omega(k\log(n/k)+1)$ is given, which shows that their first two upper bounds are asymptotically optimal, and the third is within a $O(\log\log n)$-factor of optimal.
When a global clock is available, Jurdziński and Stachowiak (jurdzinski2015cost, ) show that even with a global block, any deterministic algorithm must incur $\Omega(\log^{2}n)$ latency. This may appear to be a weak lower bound; however, $O(\log^{2}n)$ latency is possible with a deterministic algorithm under the SINR model (recall Section 3.3.2) as shown by Jurdzinski et al. (jurdzinski2013distributed, ).
@@
Discussion. Much of the theoretical territory has been mapped out with nearly-matching upper and lower asymptotic bounds on latency, and the results are less bleak than those surveyed under the AQT model. Notably, De Marco and Kowalski (DEMARCO20171, ; de2013contention, ) demonstrate that deterministic acknowledgement-based algorithms can provide latency that is close to the asymptotic optimal—differing only by an $O(\log n\log\log n)$-factor—even with without CD, a global clock, knowledge of $k$, or the time when the first station awakens.
More generally, we see several results that deviate from an asymptotically-optimal latency by (roughly) a logarithmic factor. The presence of this term may seem puzzling at first glance, but it arises from first deriving a randomized algorithm that succeeds with high probability in $n$ (which incurs the logarithmic factor), and then showing the existence of a deterministic algorithm via the probabilistic method. This technique is what is often providing the logarithmic-factor discrepancy between upper and lower bounds.
Despite this improvement, deterministic approaches fall short of being able to offer $O(k)$ latency in the dynamic setting; this is unsurprising given the analogous situation in the static setting (Section 2.2). A natural method to seek improvement is via randomized algorithms, which we survey next.
@@ 3.3. Randomized Approaches for Contention Resolution
As we saw in Section 2, randomized backoff algorithms provide asymptotically-optimal makespan for the static setting without requiring unique IDs, CD, or knowledge of $n$. In contrast to the results surveyed in Sections 3.1 and 3.2, randomized approaches seem likely to perform better in the dynamic setting.
@@
When a packet $u$ arrives, it listens to the channel to determine whether it is already in use by other packets. Any packets that have arrived ahead of $u$ may still be attempting to succeed and, if so, they should also be emitting a “busy signal”, which we describe momentarily. If $u$ detects this busy signal, then it listens until this signal ceases, at which point it can start attempting to transmit. In this way, there are always two groups of packets: (1) active packets that are trying to transmit, while maintaining a busy signal, and (2) idle packets that are waiting for their turn on the channel. Given this isolation, active packets can execute a contention-resolution algorithm for the static setting in order to succeed; afterwards, any idle packets can become active and repeat this process.
@@
The busy signal can take different forms. For example, given CD and a global clock, active packets can emit a ”busy” transmission in each even-numbered slot; see Figure 6. Any idle packet that hears this signal or a collision infers that active packets are present and keeps waiting. However, having all packets send in every second slot may not be desirable, since sending consumes energy (SINHA201714, ; feeney2001investigating, ; polastre:telos, ; wang2006realistic, ). To reduce this cost, packets may emit the busy signal probabilistically, which can greatly reduce the sending cost. Under this approach, some packets may still be active, but fail to emit a busy signal. When this happens, any idle packets will become active prematurely and interfere with the already-active packets.
@@
This interference is mitigated differently, depending on the setting. Bender et al. (BenderKPY18, ; bender:contention, ) assume CD and $O(\log n)$ control bits (recall Section 1.1.4) in each transmission on the shared channel. The busy signal occurs probabilistically every $O(1)$ slots, with each active packet sending an expected $O(1)$ times to maintain the signal. Interleaved with the busy signal are slots used to estimate the number of active packets, $m\leq n$, which allows them to execute TSTB (recall Section 2.1) starting with a window of size $\Theta(m)$. Additionally, a leader is elected from the set of active packets to broadcast how many slots remain in the execution of TSTB; this serves as a busy signal during the execution of TSTB. The leader is frequently changed so that each packet is responsible for $O(1)$ transmissions in expectation. Newly-arrived idle packets need only listen for a handful of slots to receive this information, which allows them to (i) avoid becoming active prematurely and (ii) avoid any further listening until the current execution of TSTB is finished. While the above process can fail—for example, the estimate is poor, the busy signal fails, or active packets do not succeed in the execution of TSTB—this occurs with probability at most polynomially small in the number of active packets, which allows for $O(n)$ expected latency. Additionally, the expected number of channel accesses performed per packet is $O(\log\log^{*}n)$. This low number of accesses translates is important to modern networks where devices are battery-powered, given that listening and sending on the channel consumes limited energy. We note that this approach makes use of snooping (recall Section 1.1.2), where that the number of control bits sent in a packet is $O(\log n)$.
@@ 3.3.1. Randomized $k$-Selection
We extend our discussion above by continuing to survey the results that employ randomized algorithms for the dynamic $k$-Selection problem. In the context of randomized algorithms, results focus the case where $k=n$. Much of the literature can be subdivided into the no-CD and CD cases, where the latter case captures more a broader range of model aspects, such as extreme energy efficiency, along with more exotic channel feedback and adversarial behavior.
@@
No-CD. De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ) explore randomized solutions with no-CD. When $n$ is known (or a known to within a constant multiplicative factor), they provide a non-adaptive algorithm with $O(n)$ latency, which is asymptotically optimal. For the case where $n$ is unknown, a lower bound of $\Omega(n\log n/(\log\log n)^{2})$ latency is shown to hold for any non-adaptive, randomized algorithm; thus, there is an asymptotic separation in latency between adaptive and non-adaptive randomized algorithms. Additionally, the authors provide a strongly non-adaptive, non-acknowledgment-based algorithm with $O(n\log^{2}n)$ latency; this latency can be improved to $O(n\log n/(\log\log n)^{2})$ with an acknowledgment-based algorithm, which asymptotically matches their lower bound. Finally, when $n$ is unknown, the authors also provide an adaptive algorithm that achieves $O(n)$ latency
@@
A more recent result by Bender et al. (bender2020contention, ) shows that w.h.p. constant throughput can be achieved by an adaptive algorithm with no-CD. This is comparable to the adaptive algorithm by De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ) for the case where $n$ is unknown. However, the result by De Marco and Stachowiak requires that some packets remain in the system for a time even after succeeding in order to help facilitate other transmissions; in contrast, the result by Bender et al. allows packets to terminate immediately upon success. We also note that while the results of De Marco and Stachowiak and Bender et al. are efficient in terms of the number of transmission attempts made, they still require packets to listen in a constant fraction of all slots.
@@
With CD. When CD is available, Bender et al. (bender:contention, ) achieve $O(n)$ latency using expected $O(\log\log^{*}n)$ channel accesses per packet; this is shown to be asymptotically optimal by Chang et al. (ChangJP19, ). This extremely low number of accesses translates is important to modern networks where devices are battery-powered, given that listening and sending on the channel consumes energy.
When the resource may be unavailable for $D$ slots, Bender et al. (BenderFGY19, ; bender:how, ) show that $O(n+D)$ latency can be achieved with $O(\log^{2}(n+D))$ send attempts per packet; however, packets listen in each slot.
In a setting with $\mathcal{C}$ channels, Fineman et al. (fineman:contention2, ) give an algorithm with $O(\log n/\log\mathcal{C}+(\log\log n)(\log\log\log n))$ latency for the general case, and $O(\log n/\log\mathcal{C}+(\log\log n))$ in the restricted case where at most $2$ participants become active; this second case matches the lower bound provided by Newport (newport2014radio, ).
@@
Also in the multi-channel setting, Christodoulou et al. (christodoulou2018strategic, ) consider a game theoretic setting where participants may deviate from any prescribed protocol. The authors give protocols from which no participant will unilaterally deviate and which provide probablistic latency guarantees.
@@ 3.3.2. The Wakeup Problem
Many randomized solutions focus on the wakeup problem (recall Section 3.2). An early result is by Willard (willard:loglog, ), who considers the with-CD case where $k$ out of $n$ stations each has a packet to send, for $2\leq k\leq n$. Here, Willard estabishes a lower bound on “fair” protocols—for any fixed round, each device sends packets with the same probability [^7] Note that this is a different notion of fairness than described in Section 3.1, and it referred to as “uniform” by Jurdziński and Stachowiak (jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ).—showing that there is some value of $Q$ for which the expected time to success is at least $\lfloor\lg\lg n\rfloor-c$, where $c>0$ is a constant. A randomized algorithms is given with $\lg\lg n+O(1)$ latency, for when $n$ known, but $k$ is not.
In the case of no-CD, when $n$ is known and there is a global clock, Kushilevitz and Mansour (kushilevitz1998omega, ; kushilevitz1993omega, ) prove a stronger lower bound of $\Omega(\log n)$ expected rounds in the case where all devices awake in the same round (i.e., the static case); this applies even to non-uniform algorithms. For algorithms that succeed with probability at least $1-\epsilon$, this lower bound was strengthened to $\Omega(\log n\,\log(1/\epsilon))$ by Farach-Colton et al. (farach-colton:lower-bounds, ). [^8] In multi-hop networks, for uniform/fair algorithms, this lower bound is strengthened by Farach-Colton and Mosteiro to $\Omega(\log^{2}n)$ (farach2007initializing, ; farach2015initializing, ). We note a couple things. First, this is dramatically different from the lower bound by Willard (willard:loglog, ), where CD is available. Second, we note nearly-matching upper bounds are given by Jurdziński and Stachowiak (jurdzinski2002probabilistic, ), with or without a global clock, where $O(\log n\,\log(1/\epsilon))$ latency is possible so long as at least one of the following is true: devices have unique labels, or $n$ is known.
@@
Without CD, Jurdziński and Stachowiak (jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ) present randomized algorithms with error $\epsilon$ for solving several versions of the wakeup problem: where participants are labeled or unlabeled, where there is a global clock or not, and where $n$ is known or unknown. We note that, while all deterministic results rely on the use of labels, this is an example where a randomized result uses labels; in this case, the labels are used to approximate system size. The authors are concerned with algorithms where, for any fixed round, each awake device sends with the same probability. A main finding is a near-exponential separation between the global and local clock settings when $n$ is unknown and there no labeling; the former has latency (roughly) polylogarithmic in $n$ and $1/\epsilon$, while the latter requires $\tilde{\Omega}(n)$ latency. In the cases where $n$ is known or the participants are labeled (or both), the run time is $O(\log n\log(1/\epsilon))$. For fair/uniform algorithms, the authors also show a nearly-tight lower bound of $\Omega\left(\frac{\log(n)\log(1/\epsilon)}{\log\log(n)+\log\log(1/\epsilon)}\right)$ for the case where $n$ is known and the participants are labeled.
An alternative channel model makes use of the signal-to-interference-plus-noise ratio (SINR), and this has featured in several fundamental wireless-communication problems; for details, see (moscibroda2006complexity, ; jurdzinski2012distributed, ; jurdzinski2013distributed, ; moses2020deterministic, ; jurdzinski2015cost, ; JurdzinskiR17, )). Informally, the ratio of (i) the power of an incoming transmission to (ii) the sum interference from other devices plus the floor noise, must exceed a threshold value in order for the transmission to succeed. Furthermore, in this setting, collision detection is not available. Jurdziński and Stachowiak (jurdzinski2015cost, ) consider the SINR model in the absence of a global clock. They demonstrate a randomized algorithm that runs in $O(\log^{2}n/\log\log n)$ slots, assuming participants know a polynomial upper bound on the system size.
Another result in the SINR model by Fineman et al. (fineman:contention, ) shows how to solve wakeup with $O(\log n+\log R)$ latency, where $R$ is ratio between the longest and shortest link. The authors point out that, in many practical settings, $R$ is at most a polynomial in $n$, which yields an $O(\log n)$ result. [^9] This ratio appears in other related results under the SINR model, such as interference scheduling or link scheduling (e.g., (fanghanel2011improved, )).
@@
For a single-hop radio network, Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) consider $k$ devices becoming active over time, with access to a known number of $b$ channels. There is no-CD and no global clock. Here, all stations may access all channels at once. When $k$ is known, Chlebus et al. present a randomized algorithm with error $\epsilon$ that has latency $O(k^{1/b}\ln(1/\epsilon))$.
@@
Finally, we note a recent result by Gilbert et al. (gilbert2021contention, ), that examines the implications of system-size “predictions”; that is, the algorithm is given access to the probability distribution defined over the possible values of $n$. The authors give lower and upper bounds for both deterministic and randomized solutions, for the CD and no-CD cases. Interestingly, in addition to several new bounds, this work recovers some lower bounds discussed above in this section; specifically, the $\Omega(\log n/\log\log n)$ bound by Jurdziński and Stachowiak (jurdzinski2002probabilistic, ), as well as the $\Omega(\log\log n)$ lower bound by Willard (willard:loglog, ).
Discussion. For the selection problem, we can get asymptotically optimal $O(n)$ latency, and this is possible even with a high degree of energy efficiency, or if there is a powerful jamming attack. Interestingly, this latency is possible even when no-CD is available, although the result is less robust to significant jamming. For the wakeup problem, the situation is also significantly improved by randomization. Specifically, compared to the results surveyed in Section 3.2.2, we see that randomization allows for $O(\texttt{polylog}(n))$ latency in contrast to $\omega(n)$ latency in the single-hop scenario. In comparison to the results from the AQT and deterministic approaches, it is clear that randomization offers significant improvements.
Another aspect of randomized approaches to the selection problem is the focus on the number of access attempts, which translates to energy consumption. The $O(\log\log^{*}n)$ complexity achieved by Bender (bender:contention, ) exploits the ability to send and receive messages of $\Theta(\log n)$ bits. Under the more standard ternary feedback model, the best results achieve a polylogarithmic in $n$ number of access attempts.
Finally, we note that while many these randomized results achieve asymptotically optimal latency, they are fragile in the sense that even a small amount of channel interference can cause the corresponding algorithm to fail. For example, the result of Bender et al. (bender:contention, ) estimates the number of packets contending for channel access; even a single slot with interference can cause a gross overestimate. Another instance is a result by De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ), which depends on coordination messages issued by a leader; such messages can be disrupted by interference. Tolerating interference is the subject we turn to in our next section.

### v3

@@ 3.2.1. Deterministic $k$-Selection
With regard to upper bounds, De Marco and Kowalski (doi:10.1137/140982763, ) consider the case with a global clock and no-CD, as discussed above. Subsequently, De Marco et al. (DBLP:conf/icdcs/MarcoKS19, ) examine the power of acknowledgement-based versus non-acknowledgement-based deterministic non-adaptive approaches when there is no global clock and under the no-CD setting. For unknown $k$, they give a non-acknowledgement-based algorithm with latency $O(k^{2}\log n)$, and show that this can be improved to $O(k^{2}\log n/\log k)$ with an acknowledgment-based algorithm. These results are (nearly) tight; the authors also provide lower bounds on latency of $\Omega(k^{2}/\log k)$ for both cases. This illustrates a startling contrast with the case for known $k$, where the authors give an acknowledgement-based algorithm with $O(k\log k\log n)$ latency, which is known to be nearly tight (ClementiMS03, ).
@@ 3.2.2. Deterministic Wakeup
Continuing in a model with no-CD, but with a global clock, De Marco and Kowalski (DEMARCO20171, ; de2013contention, ) consider non-adaptive solutions. Here, there is a parameter $k$, for $1\leq k\leq n$, which is an upper bound on the number of stations that may become active over time. Of these stations, $s$ is the time (measured from the first slot) at which the first station becomes active. The authors look at three settings: where $s$ is known and $k$ is unknown, when $k$ is known and $s$ is unknown, and when both $k$ and $s$ are unknown. In the first two settings, the authors give algorithms that achieve a latency of $O(k\log(n/k)+1)$, while their solution for the third scenario has a latency of $O(k\log(n)\log\log(n))$. A lower bound of $\Omega(k\log(n/k)+1)$ is given, which shows that their first two upper bounds are asymptotically optimal, and the third is within a $O(\log\log n)$-factor of optimal.
Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) consider a multi-channel scenario with no-CD and no global clock. Here, all stations may access all $b\geq 1$ channels at once, and with the same parameter $k\geq 1$ being an upper bound on the number of stations that become active over time. The authors give a non-adaptive, deterministic algorithm for the cases where $k$ is unknown that has latency $O(k\log^{1/b}k\log n)$. When $b=\Omega(\log\log n)$, a faster deterministic algorithm is possible, with delay $O((k/b)\log n\log(b\log n))$, and a lower bound of $\Omega((k/b)\log(n/k))$ is shown for deterministic algorithms. For ease of presentation, we note that the authors also consider jamming (a topic discussed further in Section 4), where each channel is jammed independently and uniformly with known probability $p$ per round, where $10<p<$ 1. Here, Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) give two deterministic algorithms that succeed with probability $1-O(1/\texttt{poly}(n))$, the first with latency $O(k\log(n)\log^{1/b}k/\log(1/p))$, and the second with latency $O((k/b)\log(n)\log(b\log(n))/\log(1/p))$ when $b=\Omega(\log(b\log n))$.
@@
Finally, under the SINR channel model, Jurdziński et al. (jurdzinski2013distributed, ) give a deterministic algorithm with $O(\log^{2}n)$ makespan when there is a global clock; however, without a global clock, the authors show a $\Omega(n)$ makespan. Complementing this result, Jurdziński and Stachowiak (jurdzinski2015cost, ) show that any algorithm with access to a global clock must incur $\Omega(\log^{2}n)$ makespan in the classical (non-SINR) channel model.
Summary. Much of the theoretical territory has been mapped out with some nearly-matching upper and lower asymptotic bounds on latency, and the results are less bleak than those surveyed under the AQT model. Notably, with no-CD and assuming a global clock, we see this for both $k$-selection (De Marco and Kowalski (doi:10.1137/140982763, )) and for wakeup (De Marco and Kowalski (DEMARCO20171, ; de2013contention, )), where the latter result depends on some limited system information.
Despite providing a rosier picture, deterministic approaches for contention resolution fall short of being able to offer asymptotically-optimal latency in the dynamic case; this is unsurprising given the analogous situation in the static case (Section 2.2). Similarly, deterministic solutions to the wakeup problem yield high latency in cases without a global clock, or where key parameters are unknown. A natural place to look for improvement—perhaps even asymptotic optimality—is the domain of randomized algorithms, which we survey next.
@@ 3.3. Randomized Approaches for Contention Resolution
As we saw in Section 2, randomized backoff algorithms provide asymptotically-optimal makespan for the static case without requiring unique IDs, CD, or knowledge of $n$. Thus, there is reason to hope that randomized approaches can also provide strong results in the dynamic case.
@@
When a packet $u$ arrives, it listens to the channel to determine whether it is already in use by other packets. Any packets that have arrived ahead of $u$ may still be attempting to succeed and, if so, they should also be emitting a “busy signal”, which we describe momentarily. If $u$ detects this busy signal, then it listens until this signal ceases, at which point it can start attempting to transmit. In this way, there are always two groups of packets: (1) active packets that are trying to transmit, while maintaining a busy signal, and (2) idle packets that are waiting for their turn on the channel. Given this isolation, active packets can execute a contention-resolution algorithm for the static case in order to succeed; afterwards, any idle packets can become active and repeat this process.
@@
The busy signal can take different forms. For example, given CD and a global clock, active packets can emit a ”busy” signal in each even-numbered slot; see Figure 6. Any idle packet that hears this signal or a collision infers that active packets are present and keeps waiting. However, having all packets send in every second slot may not be desirable, since sending consumes energy (SINHA201714, ; feeney2001investigating, ; polastre:telos, ; wang2006realistic, ). To reduce this cost, packets may emit the busy signal probabilistically, which can greatly reduce the sending cost; however, there is now a probability of error in the situation where some packets may still be active, but fail to emit a busy signal. When this happens, any idle packets will become active prematurely and interfere with the already-active packets.
@@
This interference is mitigated differently, depending on the setting. Bender et al. (BenderKPY18, ; bender:contention, ) assume CD and $O(\log n)$ control bits (recall Section 1.1.4) in each transmission on the shared channel. The busy signal occurs probabilistically every $O(1)$ slots, with each active packet sending an expected $O(1)$ times to maintain the signal. Interleaved with the busy signal are slots used to estimate the number of active packets, $m\leq n$, which allows them to execute TSTB (recall Section 2.1) starting with a window of size $\Theta(m)$. Additionally, a leader is elected from the set of active packets to broadcast how many slots remain in the execution of TSTB; this serves as a busy signal during the execution of TSTB. The leader is frequently changed so that each packet is responsible for $O(1)$ transmissions in expectation. Newly-arrived idle packets need only listen for a handful of slots to receive this information, which allows them to (i) avoid becoming active prematurely and (ii) avoid any further listening until the current execution of TSTB is finished. While the above process can fail—for example, the estimate is poor, the busy signal fails, or active packets do not succeed in the execution of TSTB—this occurs with probability at most polynomially small in the number of active packets, which allows for $O(n)$ expected latency. Additionally, the expected number of channel accesses performed per packet is $O(\log\log^{*}n)$. This low number of accesses translates is important to modern networks where devices are battery-powered, given that listening and sending on the channel consumes limited energy.
We note that while a busy signal has been used to achieve strong algorithmic results, it may be problematic in practice, since a constant fraction of the active slots are wasted on this synchronization.
@@ 3.3.1. Randomized $k$-Selection
We extend our discussion above by continuing to survey the results that employ randomized algorithms for the dynamic $k$-selection problem. In the context of randomized algorithms, results focus the case where $k=n$. We note that the jamming-resistant contention-resolution results (presented next in Section 4) clearly apply when there is no jamming, although we defer their presentation until later.
@@
De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ) explore randomized solutions with no-CD. When $n$ is known (or a known to within a constant multiplicative factor), they provide a non-adaptive algorithm with $O(n)$ latency, which is asymptotically optimal. For the case where $n$ is unknown, a lower bound of $\Omega(n\log n/(\log\log n)^{2})$ latency is shown to hold for any non-adaptive, randomized algorithm; thus, there is an asymptotic separation in latency between adaptive and non-adaptive randomized algorithms. Additionally, for $n$ unknown, the authors provide a strongly non-adaptive, non-acknowledgment-based algorithm with $O(n\log^{2}n)$ latency; this latency can be improved to $O(n\log n/(\log\log n)^{2})$ with an acknowledgment-based algorithm, which comes close to matching their lower bound. Lastly, when $n$ is unknown, the authors provide an adaptive algorithm that achieves $O(n)$ latency. Informally, this algorithm employs LE to break the dynamic problem into a sequence of static contention-resolution problems, as discussed earlier in this section.
@@
A more recent result by Bender et al. (bender2020contention, ) shows that w.h.p. constant throughput can be achieved by an adaptive algorithm with no-CD. In contrast, to the adaptive algorithm (for unknown $n$) of De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ), the result by Bender et al. (bender2020contention, ) allows packets to terminate immediately upon success.
@@
When CD is available, Bender et al. (bender:contention, ) achieve $O(n)$ latency using expected $O(\log\log^{*}n)$ channel accesses per packet; this is shown to be asymptotically optimal by Chang et al. (ChangJP19, ). As discussed in Section 1.1.3, this low number of channel accesses translates is important to modern networks where devices are battery-powered, given that listening and sending on the channel consumes energy. We highlight that LE plays a key role in keeping energy costs low.
@@
In a multi-channel setting, with and without CD, Christodoulou et al. (christodoulou2018strategic, ) consider a game theoretic setting where participants may deviate from any prescribed protocol. The authors give protocols from which no participant will unilaterally deviate and which provide probabilistic latency guarantees.
@@ 3.3.2. The Wakeup Problem
There is a known separation between static and dynamic wakeup. Specifically, recall that for static wakeup with no-CD, Kushilevitz and Mansour (kushilevitz1998omega, ) show a lower bound of $\Omega(\log n)$, and an asymptotically matching upper bound is achieved via a randomized algorithm by Bar-Yehuda et al. (bar1992time, ). However, for dynamic wakeup, Jurdzinski and Stachowiak (jurdzinski2015cost, ) prove a lower bound of $\Omega(\log^{2}n)$ in the case of no-CD.
@@
Under no-CD, Jurdziński and Stachowiak (jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ) present randomized algorithms with probability of error $\epsilon$ for solving the wakeup problem under different settings: where participants are labeled or unlabeled, where there is a global clock or not, and where $n$ is known or unknown. We note that, while deterministic approaches rely on the use of labels, this is an example of labels being used with a randomization solution; in this case, the labels are used to approximate system size. The authors are concerned with uniform algorithms (recall Section 2.3). Perhaps unsurprisingly, aspects of these wakeup algorithms are reminiscent of exponential backoff. One of the main findings is a near-exponential separation between the global and local clock settings when $n$ is unknown and there no labeling; the former can be solved with latency $O(\log(n)\log(1/\epsilon))$, while the latter requires $\Omega(n/\log n)$ latency. In the cases where $n$ is known or the participants are labeled (or both), the run time is $O(\log n\log(1/\epsilon))$. For uniform algorithms, the authors also show a nearly-tight lower bound of $\Omega\left(\frac{\log(n)\log(1/\epsilon)}{\log\log(n)+\log\log(1/\epsilon)}\right)$ in the case where $n$ is known and the participants are labeled. As stated earlier in Section 2.3, this bound improved by Farach-Colton et al. (farach-colton:lower-bounds, ) to $\Omega(\log(n)\log(1/\epsilon))$, and then later shown to hold for all randomized algorithms by Newport (newport2014radio, ; newport:radio-journal, ).
Under the SINR model (recall Section 2.3), and in the absence of a global clock and assuming participants know a polynomial upper bound on the system size, Jurdziński and Stachowiak (jurdzinski2015cost, ) give a Monte Carlo algorithm that has w.h.p. $O(\log^{2}n/\log\log n)$ latency, and a Las Vegas algorithm with expected $O(\log^{2}n/\log\log n)$ latency. We note that this is better than the deterministic algorithm in (jurdzinski2013distributed, ) of $O(\log^{2}n)$ latency (recall Section 3.2.2).
@@
For a single-hop radio network, Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) consider $k$ devices becoming active over time, with access to a known number of $b$ channels. The model used is no-CD and no global clock. Here, all stations may access all channels at once. When $k$ is known, Chlebus et al. present a randomized algorithm with error $\epsilon$ that has latency $O(k^{1/b}\ln(1/\epsilon))$.
@@
Summary. For the $k$-selection problem, a key message is that we can get asymptotically optimal $O(k)$ latency. For the wakeup problem, the situation is also significantly improved by randomization. Specifically, compared to the results surveyed in Section 3.2.2, we see that randomization often allows for $O(\texttt{polylog}(n))$ latency in contrast to $\omega(n)$ latency in the single-hop scenario; however, in the local clock setting without knowledge of key parameters, the latency remains high. That said, in comparison to the results from the AQT and deterministic approaches, it is clear that randomization can offer significant improvements.
Another aspect of randomized approaches to the selection problem is the focus on the number of access attempts, which translates to energy consumption. The $O(\log\log^{*}n)$ complexity achieved by Bender et al. (bender:contention, ) exploits the ability to send and receive messages of $\Theta(\log n)$ bits. Under the more standard ternary feedback model, the best results achieve a polylogarithmic in $n$ number of access attempts.
Finally, we note that while many of the randomized results for contention resolution achieve asymptotically-optimal latency, they are fragile in the sense that even a small amount of channel interference can cause the corresponding algorithm to fail. For example, the result of Bender et al. (bender:contention, ) estimates the number of packets contending for channel access; even a single slot with interference can cause a gross overestimate. Another instance is a result by De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ), which depends on coordination messages issued by a leader; such messages can be disrupted by interference. The algorithm by Bender et al. (bender2020contention, ) is also unable to maintain asymptotically-optimal latency in the face of jamming. Thus, robustness to malicious interference presents a challenge, and it is the subject we turn to next.